2) Find the points on the given curve where the tangent line is horizontal or vertical r3 cos (0)
Answer #1

Any) briven, r=3 loso Now, Co-ordinates; converting Polar into Cartesian gora 3.8loso 2 Y - 3.21 = 3.(rloso ) n²ty²= (m?-3m)(a) Poinb 01 Horizontal tangant ; y = 3/2 & y = -3/2) Point A Point B Point of wwthical fangents; (070; } *=3) Point on Poi